OpenStudy (anonymous):

Given Polynomial: f(x) = 6x^3-7x-2 factors of P: ? factors of Q: ? Possible rational roots: ?

Parth (parthkohli):

Use the Rational Root Theorem.\[P = -2\]\[Q = 6\]

Parth (parthkohli):

Let me give you an example: If you have \(x^2 + 10x + 25\), first list the factors of \(Q\). Here, they'd be \(\pm 1\). For \(P\), \(\pm1,\pm 5,\pm 25\). So the possible rational rots are:\[\pm \dfrac{25}{1}, \pm\dfrac{5}{1},\pm \dfrac{1}{1}\]

Parth (parthkohli):

Basically, you create all combinations of the factors of \(P\) and \(Q\) and then divide \(P\over Q\).

OpenStudy (anonymous):

Thank you this helped me out alot! Thank you!